Iron metal reacts with chlorine gas. How many grams of FeCl3 are obtained when 515g Cl2 reacts with excess Fe.
1
Expert's answer
2019-10-08T07:41:05-0400

Let's write down the equation of chemical reaction (unbalanced)


"Fe + C{l_2} \to FeC{l_3}"

and balance it (balance "Cl" at first, then "Fe")


"2Fe + 3C{l_2} \to 2FeC{l_3}"

Now we can see that for each "1" mole of "C{l_2}" we get "\frac{2}{3}" moles of "FeC{l_3}" (we divide equation by 3)

So, the number of moles of "2FeC{l_3}" will be

"{\nu _{FeC{l_3}}} = \frac{2}{3}{\nu _{C{l_2}}}"

Molar mass of "C{l_2}" (atomic weight can be found in the periodic table)


"{M_{C{l_2}}} = 2 \cdot 35.45[\frac{{\text{g}}}{{{\text{mol}}}}] \approx 70.9[\frac{{\text{g}}}{{{\text{mol}}}}]"


and "{FeC{l_3}}"


"{M_{FeC{l_3}}} = 55.85[\frac{{\text{g}}}{{{\text{mol}}}}] + 3 \cdot 35.45[\frac{{\text{g}}}{{{\text{mol}}}}] \approx 162.2[\frac{{\text{g}}}{{{\text{mol}}}}]"

And we can use "\nu = \frac{m}{M}" , thus


"{m_{FeC{l_3}}} = \frac{2}{3}{m_{C{l_2}}}\frac{{{M_{FeC{l_3}}}}}{{{M_{C{l_2}}}}}"

Let's do the calculations


"{m_{FeC{l_3}}} = \frac{2}{3}515[{\text{g}}]\frac{{162.2[\frac{{\text{g}}}{{{\text{mol}}}}]}}{{70.9[\frac{{\text{g}}}{{{\text{mol}}}}]}} \approx 785.45[{\text{g}}]"